OpenStudy (anonymous):

What is the area of a circle with a diameter of twenty-eight meters? Use twenty-two sevenths for pi. six hundred sixteen square meters six hundred thirty-eight square meters six hundred sixty square meters six hundred eighty-two square meters @Venny

OpenStudy (anonymous):

Okay you remember that the diameter=2*r right, so the r=14 \(\large\ A=\pi 14^2\)

OpenStudy (anonymous):

i got 87.92

OpenStudy (anonymous):

But remember it says (22/7) for pi this time, so it is 196*(22/7)

OpenStudy (anonymous):

d?

OpenStudy (anonymous):

My calculator says its 616

OpenStudy (anonymous):

ok